OpenStudy (aleah54):

Simplify (3xy2)(4xy)(2xy)3. can you explain how you got the answer

OpenStudy (jango_in_dtown):

Hello is it

OpenStudy (jango_in_dtown):

\[(3xy^2)(4xy)(2xy)^3\]

OpenStudy (jango_in_dtown):

???

OpenStudy (aleah54):

yes

OpenStudy (jango_in_dtown):

ok... so it will be

OpenStudy (jango_in_dtown):

\[(3xy^2)(4xy)(8x^3y^3)\]

OpenStudy (jango_in_dtown):

\[=(3*4*8)(x*x*x^3)(y^2*y*y3)=96x^5y^6\]

OpenStudy (jango_in_dtown):

since a^m* a^n=a^(m+n)

OpenStudy (aleah54):

can you help me with one more